[figure: 9.7 cm]

Electronic theory for elementary and spin excitations in high-T_{c} cuprates: interplay between magnetism and superconductivity

D. Manske, I. Eremin, and K.H. Bennemann

Institut für Theoretische Physik, Freie Universität Berlin, D-14195 Berlin, Germany

Angle-resolved photoemission spectroscopy (ARPES) and inelastic neutron scattering (INS) are powerful tools for studying the elementary and spin excitations in high-T_{c} superconductors. In particular, a kink feature at energies 50±15 meV has been observed in hole-doped cuprates like Bi_{2}Sr_{2}CaCu_{2}O_{8−x}, Pb-doped Bi_{2}Sr_{2}CuO_{6}, and La_{2−x}Sr_{x}CuO_{4}[1]. In our studies we have calculated self-consistently the pronounced momentum and frequency dependence of the self-energy Σ(k, ω) originating from scattering of the quasiparticles on spin fluctuations within a unified theory for Cooper-pairing in hole-doped high-T_{c} cuprates[2]. We find that this coupling results in a kink structure in the dispersion ω(k) which agrees well with recent ARPES experiments. Furthermore, using a two-dimensional Hubbard Hamiltonian we find that the resonance peak in the magnetic susceptibility, Imχ(q, ω), appears only in the superconducting state and it scales with T_{c}[3]. By taking into account the feedback of superconductivity on Imχ(q, ω) we obtain through our calculations that tunneling data and measurements of the optical conductivity are consistent with each other and with d_{x^{2}−y^{2}}-wave pairing.

[1] A. Lanzara et al., Nature 412, 510 (2001).

[2] D. Manske, I. Eremin, and K.H. Bennemann, Phys. Rev. Lett. 87, 177005 (2001).

[3] D. Manske, I. Eremin, and K. H. Bennemann, Phys. Rev. B 63, 054517 (2001).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : D. Manske

Address for correspondence : Institut für Theoretische Physik Freie Universität Berlin Arnimallee 14, D-14195 Berlin Germany

Email address : dmanske@physik.fu-berlin.de